OpenStudy (anonymous):

volume of cone is about 1408ft^3 r is about

OpenStudy (anonymous):

v = 1408 = 1/3 pi r^2 h i cant find a value for r if dont know h

OpenStudy (anonymous):

the height of this is 21ft

OpenStudy (anonymous):

\[r =\sqrt{[3(1408)]/(21\pi)}\]

OpenStudy (anonymous):

25.13